what is "a" in; a/0.6=3a+2?
Answer by oberobic(2304)

a/.6 = 3a+7
.
a = .6*(3a+7)
.
multiply both sides by 10 to eliminate the decimal
.
10a = 6*(3a+7)
.
10a = 18a + 42
.
-8a = 42
.
a = -(42/8) = -21/4
.
Done.